OpenStudy (dumbsearch2):

Solve the following system using the graphing method. 2x - y = 5 5x + y = 2

OpenStudy (anonymous):

Rearrange them into y=f(x) form. Can you do it from there?

OpenStudy (dumbsearch2):

Well look at the attachment... I don't quite understand, ceb105?

OpenStudy (anonymous):

That question isn't clear. However it is likely asking for where the two lines intersect. Still, you need to rearrange both into the form y=f(x), does this help or do you need more?

OpenStudy (dumbsearch2):

y = 2 x-5 ?

OpenStudy (anonymous):

y=2x-5, so to help draw the line, at x=0, y=-5 and then you can see the gradient is 2 from the coefficient of x y=2-5x and with this one you can put in x=0 and see y=2. and the gradient is negative at -5, so for every one across it goes 5 down.

OpenStudy (dumbsearch2):

I got y = 2 x-5 for the first and y = 2-5 x for the second.

OpenStudy (anonymous):

Read up bud

OpenStudy (dumbsearch2):

so the answer is -5 (x) and 2 (y)?

OpenStudy (anonymous):

No, thats not right, you need to graph it, have you done that?

OpenStudy (anonymous):

solve for y and then plot the functions on your calculator. what kind of calculator do you have?

OpenStudy (whpalmer4):

You've got the equations in standard form: Ax + By = C, A and B both non-zero. The x-intercept and y-intercept are given respectively by (C/A,0) and (0,C/B). Plot those two points and draw a line through them. Repeat with the other equation. Pick off the point of intersection. Hint: x = 1.

OpenStudy (anonymous):

2x - y = 5 (1) 5x + y = 2 (2) From 1, -y = 5 - 2x y = -5+2x From 2. 5x -5 + 2x = 2 7x - 5 = 2 7x = 7 x =1

OpenStudy (anonymous):

so y =-3

OpenStudy (whpalmer4):

Nice use of the graphical method there, not!